### v1

# A Dataset for Large Language Model-Driven AI Accelerator Generation

Abstract
In the ever-evolving landscape of Deep Neural Networks (DNN) hardware acceleration, unlocking the true potential of systolic array accelerators has long been hindered by the daunting challenges of expertise and time investment. Large Language Models (LLMs) offer a promising solution for automating code generation which is key to unlocking unprecedented efficiency and performance in various domains, including hardware descriptive code. However, the successful application of LLMs to hardware accelerator design is contingent upon the availability of specialized datasets tailored for this purpose. To bridge this gap, we introduce the Systolic Array-based Accelerator DataSet (SA-DS). SA-DS comprises of a diverse collection of spatial arrays following the standardized Berkeley’s Gemmini accelerator generator template, enabling design reuse, adaptation, and customization. SA-DS is intended to spark LLM-centred research on DNN hardware accelerator architecture. We envision that SA-DS provides a framework which will shape the course of DNN hardware acceleration research for generations to come. SA-DS is open-sourced under the permissive MIT license at this https://github.com/ACADLab/SA-DS.
Index Terms:
Systolic array design, LLM-powered hardware synthesis, accelerator architecture optimization, EDA
1 Introduction
Artificial Intelligence (AI) has shown a remarkable potential to address complex design problems ranging from software development to drug discovery. A key advantage of AI is a significant reduction of the manual effort and expertise requirements. This promising capability of AI suggests its application in hardware design, particularly for developing specialized AI accelerators needed to keep pace with the rapid evolution of Deep Neural Networks (DNNs) [1]. In hardware design for DNNs, the complexity and need for expert knowledge have been major limitations [2, 3].
Systolic array accelerators, typically obtained with specialized AI hardware generators such as Gemmini [3], have significantly advanced the processing capabilities for DNNs, providing high throughput and energy efficiency. These systems, integrated with architectures like the Rocket Chip processor [4], demonstrate the scalability and flexibility necessary for contemporary AI applications. Despite these appealing achievements, challenges such as the low-level nature, the complex programming interfaces, memory usage, and the need for extensive development times persist [5]. Moreover, systolic array accelerator generators like Gemmini [3] generally face limitations in efficiently handling diverse and irregular computational patterns beyond their optimized standard operations [6]. These limitations underscore the need for innovative solutions such as AI model-based solutions [5, 6].
At the frontier of AI advancement, Large Language Models (LLMs) [7] offer an appealing solution for alleviating the challenges in hardware accelerator design. Along this line, GPT4AIGChip [8] exemplifies using LLMs to automate the hardware design process, from conceptual design to synthesis to fabrication. However, the lack of specialized datasets of hardware accelerator design artifacts presents a strong obstacle in fully leveraging the potential of LLMs [9]. This limitation restricts usage to the vanilla LLMs without fine-tuning or in-context learning [9]; two of the most effective approaches for maximizing the LLM capabilities.
To bridge the gap, we introduce a Systolic Array Accelerator Dataset (SA-DS) to facilitate effective learning and generation of optimized designs by LLMs. Specifically, our contributions include:
(1) We create, curate, and release SA-DS, the first systolic array accelerator dataset for DNN hardware accelerator design. Each data point in SA-DS features a verbal description of an accelerator micro-architecture and a Chisel description of the design itself. These accelerator designs are obtained using the Gemmini generator [3].
(2) We demonstrate the potential of SA-DS in enabling LLM-based hardware accelerator design by showcasing its suitability for generating viable accelerator designs by providing 1-short prompts with multiple LLMs. Experimental results validate the suitability of SA-DS in providing high-quality and relevant accelerator design examples to contemporary LLMs including GPT [7], Claude [10], and Google’s Gemini [11] as compared to existing Hardware Description Language (HDL) datasets.
Figure 1: Gemmini architectural template for ASIC accelerator design [3] .
2 Background
Static Accelerator Generation Tools: A wide array of static DNN accelerator design tools have been developed such as VTA [12], MAGNet [13], and DNNWeaver [14] to suit various applications. These tools provide many hardware architecture templates supporting vector/systolic spatial arrays, data flows, software ecosystems, and OS support. Among these tools, Gemmini [3] is a comprehensive and well-packaged open-source infrastructure tailored to design full-stack DNN accelerators. Gemmini offers a versatile hardware framework, a multi-layered software stack, and an integrated System-On-Chip (SoC) environment based on the architectural template shown in Fig. 1. Central to a design provided by Gemmini is a spatial array architecture, where the template employs a 2-D array of tiles containing Processing Elements (PE). These PE units operate in parallel, handling Multiply-Accumulate (MAC) operations efficiently. To optimize the area, power, and performance trade-offs, the size of the spatial array, input type, and function units such as non-linear functions, pooling, normalization, and dataflow can be adjusted.
TABLE I: Comparison of the state-of-the-art LLM-based HDL/HLS generators.
Function/Property Ours ChatEDA [9] VeriGen [15] GPT4AIGChip [8] ChipGPT [16] Chip-Chat [17] AutoChip [18]
Function AI Accelerator Gen. RTL-to-GDSII Verilog Gen. AI Accelerator Gen. Verilog Gen. Verilog Gen. Verilog Gen.
Chatbot
Dataset NA ✓(Verilog) NA NA NA
Output format Chisel GDSII Verilog HLS Verilog Verilog Verilog
Automated Verification
Multi-shot examples
Human in Loop Low NA Medium Medium Medium High Low
A user interface featuring Prompt Optimization for the input of LLM. Not applicable.
LLM for Hardware Design: LLMs are increasingly pivotal in generating HDL and High-Level Synthesis (HLS) code. Table I provides a high-level comparison of the state-of-the-art methods along this line. GitHub Copilot [19] pioneered automatic code generation, setting the stage for domain-specific applications like DAVE [20]. Expanding on these capabilities, VeriGen [15] and ChatEDA [9] advance the field by refining hardware design workflows and automating the Register-Transfer Level (RTL) to Graphic Data System version II (GDSII) process with fine-tuned LLMs. ChipGPT [16] and Autochip [18] further this evolution by integrating LLMs to generate and optimize hardware designs, with Autochip producing accurate Verilog code through simulation feedback. Chip-Chat [17] demonstrates the efficacy of interactive LLMs like ChatGPT-4 in fast-tracking design space exploration. RTLLM [21] and GPT4AIGChip [8] specifically target design process efficiency, highlighting LLMs’ capacity to manage complex design tasks and expand access to AI accelerator design, showcasing the broad potential of LLMs in hardware design. However, except for GPT4AIGChip [8], these works do not address using LLMs for AI hardware accelerator architecture design.
3 SA-DS and An Envisioned Framework
We propose a dataset for LLM-enhanced AI hardware accelerator design and envision a general framework for its applicability. Due to space limitations, we restrict our presentation to describing how it can be utilized. This framework is outlined in Fig. 2. Since an LLM’s response is determined by both the prompt and the model coefficients, the framework focuses on these two aspects. An immediate usage of SA-DS in the envisioned framework is to help to fine-tune a generic LLM for the task of hardware accelerator design (Step 2 in Fig. 2). This can be achieved by allowing its samples to partially or fully alter the model coefficients. Besides, multi-shot prompting techniques such as in-context learning can be used where SA-DS will function as the source for multi-shot examples. Also, an initial prompt conveying the user’s intent and key software and hardware specifications of the intended design can still be further engineered/optimized through the Prompt Optimizer step (Step 1). However, administering this optimization requires the development of timely and accurate evaluation techniques and metrics for the designs generated. Since SA-DS combines verbal description and systolic array design pairs, a systolic array accelerator is taken as an outcome from the LLM (step 3). Next, a third-party quality evaluation tool can be utilized to provide a quantitative evaluation of the design, verify functional correctness, and integrate the design with the full stack. (step 4).
Figure 2: An envisioned framework for utilizing SA-DS.
In the proposed usage framework, once the LLM generates systolic array accelerators (step 3), the process moves to quality and functional evaluation (step 4). The design, often formulated in Chisel, undergoes conversion to RTL using tools like Verilator to assess functional correctness and full-stack integration, thus translating into a verifiable HDL. This step crucially embeds an automated RTL-GDSII validation process, where generated designs are assessed and flagged as Valid or Invalid based on their code sequence completeness and input-output correctness. Valid designs advance to resource validation, focusing on optimizing for Power, Performance, and Area (PPA) metrics. Conversely, designs flagged as Invalid trigger a feedback loop for error analysis and LLM retraining, facilitating iterative refinement (Steps 2 to 5) to meet the set performance criteria. Ultimately, the process culminates in Step 6, where a script generates Tool Command Language (TCL) instructions to automate RTL evaluation for HDL codes, integrating with synthesis tools to comprehensively assess and validate the PPA metrics of the hardware design.
4 SA-DS Sample Creation
SA-DS uses the Gemmini generator to provide a variety of spatial array designs, making it easier for users to adapt and reuse these designs for different projects. The dataset and design tools are developed with Chisel, a programming language embedded in Scala [22], known for its clear and efficient coding style [23]. Gemmini’s configurable nature allows for significant customization, suiting various application-specific requirements, thus supporting the advancement in AI chip design [24]. This combination of a versatile template like Gemmini and a powerful design language like Chisel ensures that SA-DS can effectively meet the diverse needs of hardware design in AI applications.
Algorithm 1 describes how SA-DS is created within the Chipyard framework [25], which ensures that the designs are verifiable. The process focuses on generating spatial array structures and function units from the Gemmini codebase. Specifically, the algorithm iterates through various configurations of these elements, as indicated in line 6, guided by insights from analyzing the Gemmini code. Each modification made during this process is checked for accuracy using Verilator, ensuring that each version of the design (M) is properly annotated to highlight its key features. The variables and their values used in this algorithm are carefully chosen based on extensive testing with the Gemmini template, leading to a diverse set of potential Gemmini designs, as shown in Fig. 3.
Algorithm 1 SA-DS Creation with the Gemmini Generator
1:Input: Source Code $S$
2:Output: List of Verified Modified Source Codes $M$
3:$P\leftarrow$ list of changeable variable parameters
4:$M\leftarrow$ empty list
5:function GenerateVariations($S,P$)
6:    for each $combination$ in $P$ do
7:         $S_{mod}\leftarrow S$
8:         for each $(parameter,value)$ in $combination$ do
9:             Replace $parameter$ in $S_{mod}$ with $value$
10:         end for
11:         $verified\leftarrow$ VerifyWithVerilator($S_{mod}$)
12:         if $verified$ then
13:             $M.append(S_{mod})$
14:         end if
15:    end for
16:    return $M$
17:end function
18:function VerifyWithVerilator($S_{mod}$)
19:    return Verilator verification result for $S_{mod}$
20:end function
The SA-DS, generated as detailed in Algorithm 1, offers a variety of configurations influenced by Function Unit (FU) availability and spatial array sizes, leading to a structured dataset easily navigable and applicable to diverse hardware design needs. As illustrated in Fig. 3, the dataset organizes these configurations into six main categories, each containing 1536 unique samples. These samples are enriched with dataflow variations like Output Stationary (OS), Weight Stationary (WS), and their combinations, accounting for 512 samples per dataflow type. The distribution of these parameters and their corresponding function units is systematically represented in Fig. 4, facilitating an understanding of the dataset’s comprehensive nature and the interplay between different function units in each configuration.
Figure 3: The design space parameters of the proposed SA-DS based on Gemmini[3].
Significance of the Parameters: M represents the configurations derived from Gemmini, maintaining full-stack compatibility. These configurations are crucial for defining the hardware accelerator’s micro-architectural elements based on Gemmini’s template, including scratchpad and accumulator sizes. Key parameters include:
  • Spatial Array Size: Defines the number of PEs, crucial for computational capacity.
  • DataFlow: Manages data movement among PEs, with options like OS, WS, or an automatic selection during runtime.
  • Function Units: Additional units that support DNN functionalities like ReLU and normalization.
  • Accumulation & Spatial Array Output Type: Affects computation precision, primarily supporting signed integer types, with potential expansion to floating-point and complex integer types.
These elements facilitate customization to meet specific application requirements.
Figure 4: The frequency of sample sin SA-DS in terms of (a) function units in each category based on the Data flow for systolic arrays, (b) Function units available individually or in combination with the others.
5 Experimental Analysis
In this section, we evaluate the effectiveness of SA-DS in supporting the design generation process for hardware accelerators via LLMs, referencing the framework depicted in Fig. 2. Due to space limitations and given the complexity of LLM fine-tuning and prompt optimization requiring further research, our analysis is conducted conceptually. We initiate a proof-of-concept experiment that benchmarks SA-DS against a recent HLS Dataset (HLSD) from [26]. This experiment considers utilizing each dataset to supply one-shot examples for LLM prompts, aiming to enhance the generation of hardware designs from verbal descriptions. To objectively assess the impact of each dataset, we analyze the code quality derived from representative prompt-code pairs selected from SA-DS. The experiment extends across four prominent LLMs; GPT-4 [7], GPT-3.5 [27], Claude [10], and Gemini Advanced [11]. Reflecting the diversity of hardware specifications covered by SA-DS, our methodology includes randomly selecting test sets from six categories within SA-DS, ensuring each category is represented by 30 samples.
Evaluation is conducted by manual code review with an HLS and Chisel expert. Due to the manual nature of verification, a bi-state verification scheme is adopted. A pass characterizes the generation of a complete and functional code complying with the verbal description or the case where the LLM generates the most crucial portions of the code and leaves redundant lines or if the code is extended to have other unended functionalities beyond what is requested in the verbal description. Conversely, a fail refers to generating incomplete code, incorrect file headers, or incurring fatal errors of different types that render the code unfunctional. Therefore, we use Verilator as an automated design verification tool exclusively for codes marked as Pass. The results of this experiment are summarized in Table II.
TABLE II: Suitability of 1-shot examples: SA-DS vs. HLSD
LLM SA-DS HLSD
Pass Fail Pass Fail
GPT-4 135 45 72 108
Gemini Advanced 144 36 57 123
GPT-3.5 155 25 68 112
Claude 150 30 71 109
The comparison between SA-DS and HLSD datasets in generating one-shot prompts for LLMs like GPT-4, Gemini Advanced, GPT-3.5, and Claude, as shown in Table II, reveals a clear pattern. SA-DS consistently shows fewer failures and more passes across all tested LLMs with around 46% more passes on average. This suggests that SA-DS’s examples better align with the LLMs’ capabilities, leading to more effective code generation. The higher Pass rates in SA-DS imply that while not perfect, the generated code often needs fewer revisions to meet design requirements, indicating its practical value in streamlining the accelerator design process.
6 Conclusion
This study has introduced the first publicly accessible LLM prompt-Chisel code dataset, dubbed SA-DS. The prompt and code examples in SA-DS cover wide variety of applications and design criteria. A proof-of-concept experiment has showcased the benefits of SA-DS in enabling the high-quality generation of hardware accelerator designs with mere verbal descriptions of novice users. This exemplifies the promising potential of enabling further research in the area of utilizing LLMs for automated hardware design generation. Key examples along this line include fine-tuning high-end LLMs for hardware design, optimized multi-shot learning, and prompt engineering serving the objectives of design efficiency in terms of execution time, hardware cost, and power consumption.